OpenStudy (anonymous):

proof if tan^-1x + tan^-1y + tan^-1z= pi/2 then xy +yz+zx=1

OpenStudy (anonymous):

this is what i done so far \[\tan ^{-1} \left( \frac{ x+y }{ 1-xy } \right) + \tan ^{-1}z =\pi/2\]

OpenStudy (anonymous):

\[\frac{ x+y }{ 1-xy } +z = \tan \pi/2\]

OpenStudy (anonymous):

let a = arctan(x) b = arctan(y) c= arctan(z) then x = tan a y = tan b z = tan c xy + yz + zx = (tan a)*tan(b) + (tan b ) * tan c + tan c * tan a

OpenStudy (anonymous):

forgot to mention i need to use \[\tan ^{-1}x + \tan ^{-1}y= \tan ^{-1}\left( \frac{ x+y }{ 1-xy } \right)\]

Parth (parthkohli):

Let \(\tan^{-1}x=\alpha, \tan^{-1}y = \beta, \tan^{-1}y = \gamma\).\[\alpha + \beta + \gamma =\pi/2\]\[\tan (\alpha + \beta + \gamma) = \frac{\tan \alpha + \tan \beta + \tan \gamma - \tan \alpha \tan \beta \tan \gamma}{ 1 - \tan \alpha \tan \beta - \tan \beta \tan \gamma - \tan \gamma \tan \alpha }\]\[= \frac{x+y+z-xyz}{1-(xy+yz+zx)}=\tan \pi/2\]In order to be undefined, \(xy+yz+zx = 1\).

OpenStudy (irishboy123):

from what you've done so far \(\tan ^{-1} \left( \frac{ x+y }{ 1-xy } \right) + \tan ^{-1}z =\pi/2\) \(\tan ^{-1} \left( \frac{ x+y }{ 1-xy } \right) =\pi/2 - \tan ^{-1}z\) then take tans on both sides....

OpenStudy (anonymous):

@IrishBoy123 done

Parth (parthkohli):

You can memorise the formula:\[\tan (x_1 + x_2 + x_3 + \cdots) = \frac{S_1 - S_3 + S_5 - S_7 + \cdots}{S_0 - S_2 + S_4 - S_6 + \cdots}\]Where \(S_n\) is the cyclic sum of tangents taken \(n\) at a time.

OpenStudy (loser66):

Apply the given information twice, you can get the answer \(tan^{-1}x + tan^{-1}y= tan^{-1}\dfrac{x+y}{1-xy}\) now, \(\dfrac{x+y}{1-xy}is X, ~and~z~is~Y\) You have \(tan^{-1}X+tan^{-1}Y= tan^{-1}\dfrac{X+Y}{1-XY}=\dfrac{\pi}{2}\) That is \(tan (\pi/2) = \dfrac{X+Y}{1-XY}\) and we know that tan (pi/2) is undefined, hence \(1-XY=0\implies XY=1\) As above, \(X =\dfrac{x+y}{1-xy}, Y =z\), hence \(XY = \dfrac{x+y}{1-xy}z=1\) that is (x+y)z = 1-xy, hence xz +yz +xy =1

OpenStudy (anonymous):

@Loser66, when pi/2 is undefined, what happen to the numerator, X+Y?

OpenStudy (loser66):

a fraction is undefined when the denominator =0 no matter what the numerator is. So that we don't care about the numerator.

OpenStudy (loser66):

Note: pi/2 is not undefined, but tan (pi/2) is undefined.

OpenStudy (anonymous):

thank you so much, i finally understand

OpenStudy (anonymous):

$$ \large { \tan ^{-1} \left( \frac{ x+y }{ 1-xy } \right) + \tan ^{-1}z =\pi/2 \\ \tan ^{-1} \left( \frac{ x+y }{ 1-xy } \right) =\pi/2 - \tan ^{-1}z \\ \tan(\tan ^{-1} \left( \frac{ x+y }{ 1-xy } \right) )=\tan(\pi/2 - \tan ^{-1}z) \\ \frac{ x+y }{ 1-xy }=\cot( \tan ^{-1}z) \\ \frac{ x+y }{ 1-xy } =1 / (\tan( \tan ^{-1}z)) \\ \frac{ x+y }{ 1-xy } =1 / z }$$

OpenStudy (anonymous):

@jayzdd, thanks :)

OpenStudy (anonymous):

credit goes to Parthkholi + irishboy

OpenStudy (anonymous):

thank you so much guys, could not done without you

OpenStudy (anonymous):

@jayzdd why there is cot??

OpenStudy (anonymous):